OpenStudy (anonymous):

a wire 34 cm long is cut into 2 pieces. One piece is bent to form a square and the other is bent to form a rectangle that is twice as wide as it is long. How should the wire be cut in order to minimize the total area of the square and the rectangle?

OpenStudy (anonymous):

You're given that the total perimeter of both wire shapes is given by \[34=4x+2y+2z,\text{ where}\] x = length of side of square, y = length of rectangle, and z = width of rectangle. You're told that the rectangle's width is twice its length, so you know \(z=2y.\) Substituting, your perimeter equation becomes \[34=4x+2y+4y\\ 34 = 4x + 6y\] The total area of both pieces is given by \[A=x^2+yz,\text{ or, using the sub from earlier,}\\ A=x^2+2y^2\] Find an expression for x in terms of y (or y in terms of x) which you will sub into the area equation. From there, you apply the derivative test and find the y (or x, depending on your expression) that minimizes the total area.

OpenStudy (anonymous):

thank you so much! but can you help me solve it? I dont know how to find the expression in terms of x or y and then apply the derivative test. Thank you!

OpenStudy (anonymous):

\[4x+6y=34\Rightarrow \color{red}{y=\frac{17-2x}{3}}\] \[\begin{align*}A&=x^2+2y^2\\ &=x^2+2\left(\color{red}{\frac{17-2x}{3}}\right)^2\\ &=x^2+2\left(\frac{289-68x+4x^2}{9}\right)\\ A&=\frac{17}{9}x^2-\frac{136}{9}x+\frac{578}{9}\end{align*}\] Now apply the first derivative test, i.e. find the critical points of A (values of x that make A' = 0), then determine over which intervals A is increasing or decreasing. \[A'=\frac{34}{9}x-\frac{136}{9}\] Set \(A'=0:\) \[0=\frac{34}{9}x-\frac{136}{9}\Rightarrow x=4\] The intervals you thus have to check are \((0,4)\text{ and }(4,\infty)\). This is because x can't be negative or zero (if x = 0, you don't have a square, which means you don't have the two pieces of wire to begin with). Over the first interval, pick a convenient x-value (such as \(x=1\)), and evaluate the derivative for that value. Sticking with 1, you have \(\displaystyle A'(1)=-\frac{102}{9}\), which tells you that A is decreasing. (Keep in mind that \(A'>0\Rightarrow A\text{ increasing, and }A'<0\Rightarrow A\text{ decreasing.} \)) Over the second interval, pick \(x=5\) and find that \(\displaystyle A'(5)=\frac{34}{9}>0\Rightarrow A\text{ increasing.}\) The decrease-then-increase indicates a minimum, so the test confirms that a minimum occurs for x = 4.

OpenStudy (anonymous):

Since the minimal area occurs for x = 4, you'd cut a piece of wire with length 4*4=16 for the square piece, and leave the rest for the rectangular piece.